# E4S: Fine-grained Face Swapping via Editing With Regional GAN Inversion

### v2

@@ Abstract
This paper proposes a novel approach to face swapping from the perspective of fine-grained facial editing, dubbed “editing for swapping” (E4S). The traditional face swapping methods rely on global feature extraction and often fail to preserve the source identity. In contrast, our framework proposes a Regional GAN Inversion (RGI) method, which allows the explicit disentanglement of shape and texture. Specifically, our E4S performs face swapping in the latent space of a pretrained StyleGAN, where a multi-scale mask-guided encoder is applied to project the texture of each facial component into regional style codes and a mask-guided injection module then manipulates feature maps with the style codes. Based on this disentanglement, face swapping can be simplified as style and mask swapping. Besides, since reconstructing the source face in the target image may lead to disharmony lighting, we propose to train a re-coloring network to make the swapped face maintain the lighting condition on the target face. Further, to deal with the potential mismatch area during mask exchange, we designed a face inpainting network as post-processing. The extensive comparisons with state-of-the-art methods demonstrate that our E4S outperforms existing methods in preserving texture, shape, and lighting. Our implementation is available at https://github.com/e4s2023/E4S2023.
@@
Figure 1: Compared with the existing StyleGAN-based face swapping approaches MegaFS [1], StyleFusion [2], and HiRes [3], our E4S can achieve high-fidelity swappped results. Note that ours achieves best identity preservation and proper occlusion handling, where the target skin tone is kept and the target lighting condition is well transferred to the swapped results. All the facial images are at $1024\times 1024$.
1 Introduction
@@
Face swapping aims at not only transferring the identity information (e.g., shape and texture of facial components) of a source face to a given target face, but also keeping the identity-irrelevant attribute information unchanged (e.g., expression, head pose, lighting, and background, etc.). It has broad masses of potential applications in movie composition, computer games, and virtual human broadcasting, thus attracting considerable attention in the field of computer vision and graphics.
Early face-swapping works [4, 5] attempt at pixel-level manipulation for exchanging inner face regions, sometimes with the help of 3D facial models to handle the pose difference between the two input faces [6, 7]. These methods typically have difficulties in preserving the target facial attributes (e.g., expression) [4] and the source identity [7]. Recent methods [8, 9, 1, 10] employ Generative Adversarial Networks (GAN) [11] to achieve high fidelity swapping results. Although continuous efforts have been pursued, there are two long-standing challenges in face swapping that remain unsolved, as we will elaborate next.
@@
The first challenge is identity preservation, i.e., making the swapped results faithfully preserve the facial characteristics of the source person. Most mainstream approaches  [8, 9, 10] extract the global identity-related features and inject them into the face generation process, where these identity features are obtained via a pre-trained 2D face recognition network [12] or a 3D morphable face model (3DMM) [13, 14]. Nonetheless, these methods always exhibit the ”in-between effect” and struggle to present satisfying results. Concretely, their swapped result looks like a third person, which resembles both the source and target faces. Here, we argue the potential reason is current identity extractors are mainly designed for classification rather than generation, thus cannot capture some informative and important facial visual details. Moreover, when a 3D face model takes one single image as input, it can hardly achieve precise shape recovery robustly.
@@
The second challenge is how to deal with facial occlusion. It is common that some face regions are occluded by hair (or eyeglasses) in the input images. An ideal swapped result should maintain the hair (or eyeglasses) of the target. That is, it needs to generate the occluded facial regions for the source face. To deal with occlusion, FSGAN [15] trains an inpainting sub-network to generate the missing pixels of the source. However, their inpainted faces are blurry. FaceShifter [9] designs a refinement network to recover the occluded region in the target. Nonetheless, the refinement network always brings back certain identity information of the target to the source image.
@@
Although past efforts have achieved promising results, the above challenges still pose the main obstacles that prevent fine-grained face swapping. Here comes to a question that can we build a face swapping framework to handle the aforementioned challenges at once? In this paper, we will explore such feasibility. Specifically, we propose to rethink face swapping from a new perspective of fine-grained face editing, i.e., “editing for swapping” (E4S). The key insight behind this is that given the disentangled shape and texture of individual facial components from two faces, face swapping can be transformed into the problem of local shape and texture replacement between them.
@@
Inspired by the previous fine-grained face editing approach [16], we apply separated component masks for local feature extraction. Then, we recompose local shapes and textures, which are fed into a mask-guided generator to synthesize the swapped result. Before conducting shape and texture swapping, we use a face reenactment model [17] to drive the source person to have the same pose and expression as the target. The unique advantage of our E4S framework is that the facial occlusion challenge can be naturally addressed by facial masks since the face parsing network [18] can provide labels of each face region. Our mask-guided generator can fill out the swapped masks with the swapped texture features adaptively. It does not rely on additional dedicated modules as in previous efforts [9, 15].
@@
Given that StyleGAN [19] has achieved striking performance on high-quality image editing, the core of our E4S framework is taking advantage of a pre-trained StyleGAN [19] for the disentanglement of shape and texture. This is a demanding mission, since the existing GAN inversion methods [20, 21, 22] only performs global attribute editing (e.g., age, gender, and expression) in the global style space of StyleGAN, and does not throw any glance on local shape and texture editing. To achieve this goal, we propose a novel Regional GAN Inversion (RGI) method that incorporates facial masks into style embedding and introduces a regional $\mathcal{W}^{+}$ space, indicated as $\mathcal{W}^{r+}$. Specifically, we design a mask-guided multi-scale encoder that maps an input face into the style space of StyleGAN, where we make each facial component have a set of style codes for different layers of the generator. Besides, based on given facial masks, a mask-guided injection module is proposed, which uses style codes to manipulate the feature maps in the generator. That is, the disentanglement of texture and shape is conducted by using style codes and masks to represent them, respectively. With such disentanglement, fine-grained face swapping can be achieved by flexibly swapping the style codes and masks of all facial components.
During the mask-exchange process, it is difficult to obtain a perfect recomposed mask for the swapped face, especially when the face shapes of source and target are very different. For example, when source’s face is thin and target’s face is fat, some mismatched facial mask regions need to be replaced by background, neck, or hair, etc. In this paper, we propose to train a face inpainting network as post-processing, where we form the paired data by erasing the mismatched area from the source as the before-inpainting data, and using the original source as the after-inpainting data. During inference, guided by those mismatched mask regions, the inpainting network can generate swapped results which meet the face shape of the source.
@@
Besides shape and texture, we argue that lighting is also crucial for face swapping. If the lighting condition of the source and target is extremely different, it will be unnatural to combine the source face and the surrounding ambient from the target. To deal with this, inspired by Heser [23], we reformulate the lighting-transferring problem as conducting face re-coloring from its corresponding grayscale images. Concretely, we train a face re-coloring network in a self-supervised manner. It is trained by paired data with the same identity, where one is a grayscale face while the other is the corresponding colored face. Note that skin tone and lighting are entangled, the results of our face re-coloring network would maintain the skin of the target, nonetheless, which is more practical in real face swapping application. In summary, the main contributions of this paper are:
@@
  • We propose E4S, a fine-grained face swapping framework from a new perspective of face editing. Our high-fidelity face swapping is capable of preserving source identity and target lighting and dealing with occlusion challenge.  [^1] The implementation is at https://github.com/e4s2023/E4S2023. The project page is available at https://e4s2023.github.io/.
@@
  • we train a face re-coloring network, making the lighting of the swapped result be faithful to the target.
  • To solve the mismatch area that may occur in the process of mask exchange, we additionally designed a face inpainting network for post processing for E4S.
  • Extensive quantitative and qualitative results demonstrate that our E4s and RGI exceeds current state-of-the-art face swapping methods.
2 Related Work
GAN Inversion. There have been various approaches proposed for GAN inversion, which aim to map an image to its corresponding latent code that can faithfully reconstruct the input. Such methods are useful for image editing, as the inverted code can be modified and then fed into the generator to produce the desired output. Existing approaches for StyleGAN inversion can be broadly categorized into three groups: learning-based [21, 24, 25, 22, 26, 27], optimization-based [28, 29, 30, 31, 32], and hybrid methods [33]. Learning-based methods involve training an encoder to map the image to the latent space, whereas optimization-based methods directly optimize the latent code to minimize the reconstruction error. Although optimization-based methods tend to yield better results, they are more computationally expensive than learning-based methods. Hybrid methods seek to balance these two approaches by using the inverted code as a starting point for further optimization.
@@ 2 Related Work
Face Swapping. Face-swapping approaches can be broadly categorized into two classes: source-oriented and target-oriented [8]. Source-oriented methods [34, 35, 36, 15, 37] initiate from the source, aiming to transfer the target’s attributes to it. Early techniques in this category can be traced back to [34], which estimated 3D shape and relevant scene parameters for pose and lighting alignment. Later, [36] claimed that 3D shape estimation was unnecessary and applied face segmentation for face swapping. Recently, FSGAN [37, 15] employed a two-stage pipeline where a reenactment and inpainting network addressed pose alignment and occlusion issues, respectively. As contrast, target-oriented methods [38, 39, 8, 9, 10, 40, 41] start with the target, intending to import the source identity. Generally, these approaches maintain the source identity using a pre-trained face recognition model or 3DMMs. However, as the recognition model is trained for classification and 3DMMs lack accuracy and robustness, these methods fail to fully capture identity-related details for generation, resulting in the ”in-between effect”.
Regarding StyleGAN-based face swapping, MegaFS [1] leverages prior knowledge from the pre-trained StyleGAN, increasing image resolution to $1024^{2}$. StyleFusion [2] performs latent fusion within the $\mathcal{S}$ space [42, 43], allowing for controllable generation of local semantic regions. Nonetheless, the shape and texture of each facial region remain entangled in the $\mathcal{S}$ space. In contrast, [40] proposes a region-aware projector for adaptively transferring source identity to the target face. HiRes[3] utilizes an additional encoder-decoder for multi-scale target feature aggregation, but these two techniques do not support fine-grained and selective swapping.
@@
Figure 2: E4S framework overview. (a) We first crop the face region of the source $S$ and the target $T$ to obtain $I_{s}$ and $I_{t}$. Then, a reenactment network $G_{r}$ encourages $I_{s}$ to have a similar pose and expression towards $I_{t}$, obtaining the driven image $I_{d}$. The segmentation masks of $I_{t}$ and $I_{d}$ are also estimated. (b) Then, the driven and target pairs $(I_{d},M_{d})$ and $(I_{t},M_{t})$ are fed into the mask-guided encoder $F_{\phi}$ to extract the per-region style codes to depict the texture respectively, producing texture codes $S_{d}$ and $S_{t}$. Next, we exchange the masks and the corresponding texture codes, and send them to the pre-trained StyleGAN generator $G_{\theta}$ with a mask-guided injection module to synthesize the naive swapped face $\tilde{I}$. (c) Finally, we conduct post-processing, which includes a face re-coloring network $B_{\psi}$ for transfering the target lighting to $\tilde{I}$, multi-bend blending for pasting back the swapped face to $\rm{T}$, and a face inpainting network $P_{\tau}$ for preserving an inconsistent shape with source face.
3 Method
@@
In this section, we first introduce our proposed editing-for-swapping (E4S) framework and elaborate on each inside step in Sec. 3.1. Then, we explain the proposed region GAN inversion (RGI) method for the disentanglement of the shape and texture of facial components in Sec. 3.2. Each core module of the RGI is detailed subsequently. Next, we introduce the face recoloring network $B_{\psi}$ for maintaining the target skin tone and lighting, and the face inpainting network $P_{\tau}$ for the unnatural results caused by the possible shape mismatch during mask exchange in Sec. 3.3. After that, we present the loss functions for training in Sec. 3.4.
3.1 Editing For Swapping (E4S) framework
Given an arbitrary pair of source and target faces, our goal is to generate a swapped face that retains the identity information from the source and keeps the identity-irrelevant attributes (e.g., pose, expression, lighting and background) from the target. The pipeline of our face swapping framework is illustrated in Fig. 2, which mainly consists of three phases inside: (a) cropping and reenactment, (b) swapping and generation, and (c) post-processing.
3.1.1 Cropping and Reenactment
We first crop the face region of the source image $S$ and target image $T$ respectively, obtaining the cropped faces $I_{\rm{s}}$ and $I_{\rm{t}}$, which can be expressed as:
$I_{\rm{s}}=f_{c}(S),\quad I_{\rm{t}}=f_{c}(T),$ (1)
@@ 3.1.1 Cropping and Reenactment
where $f_{c}$ denotes the cropping processing. We use the dlib [^2] http://dlib.net/ toolbox to crop the face region and detect the facial landmarks. Then, we follow the previous method [47] to align the cropped face and resize it into 1024$\times$1024 resolution.
@@
To drive $I_{\rm{s}}$ to obtain the similar pose and expression as $I_{\rm{t}}$, we employ a pre-trained face reenactment model FaceVid2Vid [17], obtaining a driven face $I_{d}$. The face reenactment processing can be expressed as:
$I_{d}=G_{r}(I_{\rm{s}},I_{\rm{t}}),$ (2)
@@
where $G_{r}$ is the FaceVid2Vid model. Then, the segmentation masks $M_{\rm{t}}$ of the target face $I_{\rm{t}}$ and $M_{\rm{d}}$ of the driven face $I_{d}$ are estimated by an off-the-shell face parser [48]. In this way, the target and driven pairs ($I_{\rm{t}},M_{\rm{t}}$) and ($I_{d},M_{\rm{d}}$) can be formed, where each segmentation mask belongs to one of the 19 semantic categories. For brevity, we aggregate the categories of symmetric facial components, bringing 12 categories totally, i.e., background, eyebrows, eyes, nose, mouth, lips, face skin, neck, hair, ears, eyeglass, and ear rings.

### v3

@@ Abstract
This paper proposes a novel approach to face swapping from the perspective of fine-grained facial editing, dubbed “editing for swapping” (E4S). The traditional face swapping methods rely on global feature extraction and fail to preserve the detailed source identity. In contrast, we propose a Regional GAN Inversion (RGI) method, which allows the explicit disentanglement of shape and texture. Specifically, our E4S performs face swapping in the latent space of a pretrained StyleGAN, where a multi-scale mask-guided encoder is applied to project the texture of each facial component into regional style codes and a mask-guided injection module manipulating feature maps with the style codes. Based on this disentanglement, face swapping can be simplified as style and mask swapping. Besides, due to the large lighting condition gap, transferring the source skin into the target image may lead to disharmony lighting. We propose a re-coloring network to make the swapped face maintain the target lighting condition while preserving the source skin. Further, to deal with the potential mismatch areas during mask exchange, we design a face inpainting module to refine the face shape. The extensive comparisons with state-of-the-art methods demonstrate that our E4S outperforms existing methods in preserving texture, shape, and lighting. Our implementation is available at https://github.com/e4s2024/E4S2024.
Index Terms:
Face swapping, face editing, regional GAN inversion, face recoloring, face inpainting.
@@
Figure 1: Compared with the existing StyleGAN-based face swapping approaches MegaFS [1], StyleFusion [2], and BlendFace [3], our E4S can achieve high-fidelity swappped results. Note that ours achieves the best source identity and target attributes preservation, where the source skin tone and face shape are kept and the target lighting condition is well transferred to the swapped results. All the facial images except those of BlendFace [3] are at $1024\times 1024$.
footnotetext: * Eual contribution, $\dagger$ Corresponding authors.
I Introduction
@@
Face swapping aims at not only transferring the identity information (e.g., shape and texture of facial components) of a source face to a given target face, but also keeping the identity-irrelevant attribute information unchanged (e.g., expression, head pose, lighting, and background). It has broad masses of applications in movie composition, computer games, and virtual human broadcasting, attracting considerable attention in the field of computer vision and graphics.
@@
There are two long-standing challenges in face swapping field. The first is identity preservation, i.e., making the swapped results faithfully preserve the facial characteristics of the source person. Most mainstream approaches  [4, 5, 6, 7, 8] extract the global identity-related features and inject them into the face generation process, where these identity features are obtained via a pre-trained 2D face recognition network [9] or a 3D morphable face model (3DMM) [10, 11]. Nonetheless, their swapped results look like a third person, which resemble both the source and target faces. Here, we argue the potential reason is current identity extractors are mainly designed for classification rather than generation [3], thus cannot capture some informative and important facial visual details. Moreover, when a 3D face model takes one single image as input [12], it can hardly achieve precise shape recovery robustly.
@@
The second challenge is how to deal with facial occlusion. It is common that some face regions are occluded by hair (or eyeglasses) in the input images. An ideal swapped result should maintain the hair (or eyeglasses) of the target. That is, it needs to generate the occluded facial regions for the source face. To deal with occlusion, FSGAN [13] trains an inpainting sub-network to generate the missing pixels of the source. However, their inpainted faces are blurry. FaceShifter [5] designs a refinement network to recover the occluded region in the target. Nonetheless, the refinement network tends to bring back certain identity information of the target to the source.
@@
Although past efforts have achieved promising results, the above challenges still pose the main obstacles that prevent fine-grained face swapping. Here comes to a question that can we build a face swapping framework to handle the aforementioned challenges at once? In this paper, we will explore such feasibility. We propose to rethink face swapping from a new perspective of fine-grained face editing, i.e., “editing for swapping” (E4S). The key insight behind this is that given the disentangled shape and texture of individual facial components from two faces, face swapping can be transformed into the problem of local shape and texture replacement between them.
@@
Inspired by the previous fine-grained face editing approach [14], we apply separated component masks for local feature extraction. Then, we recompose local shape and texture features, which are fed into a mask-guided generator to synthesize the swapped result. Before conducting shape and texture swapping, we use a face reenactment model [15] to drive the source person to have the same pose and expression as the target. The unique advantage of our E4S framework is that the facial occlusion challenge can be naturally addressed by facial masks since the face parsing network [16] can provide labels of each face region. Our mask-guided generator can fill out the swapped masks with the swapped texture features adaptively. It does not rely on additional dedicated modules as in previous efforts [5, 13].
@@
Given that StyleGAN [17] has achieved striking performance on high-quality image editing, the core of our E4S framework takes advantage of a pre-trained StyleGAN [17] for the disentanglement of shape and texture. However, the existing GAN inversion methods [18, 19, 20] only performs global attribute editing (e.g., age, gender, and expression) in the global style space of StyleGAN, while the local control has not yet been explored. To achieve this goal, we propose a novel Regional GAN Inversion (RGI) method that incorporates facial masks into style embedding and introduces a regional $\mathcal{W}^{+}$ space, indicated as $\mathcal{W}^{r+}$. Specifically, we design a mask-guided multi-scale encoder that maps an input face into the style space of StyleGAN, where we make each facial component have a set of style codes for different layers of the generator. Besides, based on given facial masks, we propose a mask-guided injection module using style codes to manipulate the feature maps in the generator. In a word, we use style codes and masks to represent the texture and shape and conduct their disentanglement.
@@
Besides shape and texture, we argue that lighting is also crucial for face swapping. If the lighting condition of the source and target is extremely different, directly transferring the source facial color space into the target image will produce an unnatural result. To deal with this, we split the lighting transferring problem into two steps. First, we inject the source skin texture along with its lighting into the target face to obtain a naive swapped face. Second, we transfer the target lighting to the swapped face by conducting face re-coloring under the guidance of the target color space. Concretely, we train a face re-coloring network in a self-supervised manner. It is trained by paired data with the same identity, where one is a grayscale face while the other is the corresponding colored face. This two-step scheme preserves the source skin details while yielding a natural lighting of the swapped images.
The remaining problem is how to obtain a source-consistent swapped face shape since it is difficult to recompose a perfect swapped mask during exchanging masks, especially when the face shape of source and target have a large gap. To tackle this, we propose a face inpainting network to refine the face shape, which is trained through an unsupervised manner. We synthesize the paired data by erasing random areas from the original faces and train the network to predict the erased parts through visible ones based on a mismatch mask. Additionally, based on the modulation convolution [21], we propose to use the area ratio of mismatch regions to adaptively control the generation process. During inference, guided by those mismatched mask regions, the inpainting network can refine the face shape to make it preserve the shape of the source face. In summary, the main contributions of this paper are:
@@
  • We propose E4S, a fine-grained face swapping framework from a new perspective of face editing. Our high-fidelity face swapping can preserve source identity and target lighting and dealing with occlusion challenge. The implementation is at https://github.com/e4s2024/E4S2024. The project page is available at https://e4s2024.github.io/.
@@
  • We solve the lighting transferring problem by splitting it into two parts. In the first part, we integrate the source skin along with the source lighting into the target image. In the second part, we train a face re-coloring network in an unsupervised manner, which re-colors the swapped image guided by the target color space.
  • We additionally designed a face inpainting network modulated by the mismatch region ratio, which keeps the swapped face shape consistent with the source one.
II Related Work
GAN Inversion. GAN Inversion aims to map an image to its corresponding latent code that can faithfully reconstruct the input, which is useful for image editing since the inverted code can be modified and then fed into the generator to produce the desired output. Existing approaches for StyleGAN inversion can be broadly categorized into three groups: learning-based [19, 22, 23, 20, 24, 25], optimization-based [26, 27, 28, 29, 30], and hybrid methods [31]. Learning-based methods involve training an encoder to map the image to the latent space, whereas optimization-based methods directly optimize the latent code to minimize the reconstruction error. Although optimization-based methods tend to yield better results, they are more computationally expensive than learning-based methods. Hybrid methods seek to balance these two approaches by using the inverted code as a starting point for further optimization.
@@ II Related Work
Face Swapping. Face-swapping approaches can be broadly categorized into two classes: source-oriented and target-oriented [4]. Source-oriented methods [32, 33, 34, 13, 35] initiate from the source, aiming to transfer the target’s attributes to it. Early techniques in this category can be traced back to [32], which estimated 3D shape and relevant scene parameters for pose and lighting alignment. Later, [34] claimed that 3D shape estimation was unnecessary and applied face segmentation for face swapping. Recently, FSGAN [35, 13] employed a two-stage pipeline where a reenactment and inpainting network addressed pose alignment and occlusion issues, respectively. As contrast, target-oriented methods [36, 37, 4, 5, 6, 38, 39, 40, 41] start with the target, intending to import the source identity. Generally, these approaches maintain the source identity using a pre-trained face recognition model [3] or 3DMMs [12]. However, as the recognition model is trained for classification and 3DMMs lack accuracy and robustness, these methods fail to fully capture identity-related details for generation.
For StyleGAN-based face swapping, [1, 7, 8, 42] leverages prior knowledge from the pre-trained StyleGAN, increasing image resolution to $1024^{2}$. StyleFusion [2] performs latent fusion within the $\mathcal{S}$ space [43, 44], allowing for controllable generation of local semantic regions. Nonetheless, the shape and texture of each facial region remain entangled in the $\mathcal{S}$ space. Besides, [38] proposes a region-aware projector for adaptively transferring source identity to the target face. HiRes[45] utilizes an additional encoder-decoder for multi-scale target feature aggregation. However, these two techniques do not support fine-grained and selective swapping.
@@
Figure 2: E4S framework overview. (a) We first crop the face region of the source $S$ and the target $T$ to obtain $I_{\rm{s}}$ and $I_{\rm{t}}$. Then, a reenactment network $G_{\rm{r}}$ encourages $I_{\rm{s}}$ to have a similar pose and expression towards $I_{\rm{t}}$, obtaining the driven image $I_{\rm{d}}$. The segmentation masks of $I_{\rm{t}}$ and $I_{\rm{d}}$ are also estimated. (b) Then, the driven and target pairs $(I_{\rm{d}},M_{\rm{d}})$ and $(I_{\rm{t}},M_{\rm{t}})$ are fed into the mask-guided encoder $F_{\phi}$ to extract the per-region style codes to depict the texture respectively, producing texture codes $S_{\rm{d}}$ and $S_{\rm{t}}$. Next, we exchange the masks and the corresponding texture codes, obtaining $S_{\rm{swap}}$ which is then sent to the pre-trained StyleGAN generator $G_{\theta}$ with a mask-guided injection module to synthesize the naive swapped face $\tilde{I}$. (c) Finally, we propose a refinement stage, which includes a face re-coloring network $B_{\psi}$ for transfering the target lighting to $\tilde{I}$, and a face inpainting network $P_{\tau}$ for preserving a consistent shape with source face.
III Method
@@
In this section, we first introduce our proposed editing-for-swapping (E4S) framework and elaborate on each inside step in Sec. III-A. Then, we explain the proposed region GAN inversion (RGI) method for the disentanglement of the shape and texture of facial components in Sec. III-B. Each core module of the RGI is detailed subsequently. Next, we introduce the face recoloring network $B_{\psi}$ for maintaining the target skin tone and lighting, and the face inpainting network $P_{\tau}$ for the unnatural results caused by the possible shape mismatch during mask exchange in Sec. III-C. After that, we present the loss functions for training in Sec. III-D.
III-A Editing For Swapping (E4S) framework
The pipeline of our face swapping framework is illustrated in Fig. 2, which mainly consists of three phases inside: (a) cropping and reenactment, (b) swapping and generation, and (c) lighting transfer and face shape inpainting.
III-A1 Cropping and Reenactment
@@
We first use the dlib toolbox [49] to crop the face region of the source image $S$ and target image $T$ respectively, obtaining the cropped faces $I_{\rm{s}}$ and $I_{\rm{t}}$. Then, we follow the previous method [50] to align the cropped face and resize it into 1024$\times$1024 resolution.
@@
To drive $I_{\rm{s}}$ to obtain the similar pose and expression as $I_{\rm{t}}$, we employ a pre-trained face reenactment model FaceVid2Vid [15], obtaining a driven face $I_{\rm{d}}$. The face reenactment processing can be expressed as:
$I_{\rm{d}}=G_{r}(I_{\rm{s}},I_{\rm{t}}),$ (1)
@@
where $G_{r}$ is the FaceVid2Vid model. Then, the segmentation masks $M_{\rm{t}}$ of the target face $I_{\rm{t}}$ and $M_{\rm{d}}$ of the driven face $I_{\rm{d}}$ are estimated by an off-the-shell face parser [51]. In this way, the target and driven pairs ($I_{\rm{t}},M_{\rm{t}}$) and ($I_{\rm{d}},M_{\rm{d}}$) can be formed, where each segmentation mask belongs to one of the 19 semantic categories. For brevity, we aggregate the categories of symmetric facial components, bringing 12 categories totally, i.e., background, eyebrows, eyes, nose, mouth, lips, face skin, neck, hair, ears, eyeglass, and ear rings.
III-A2 Swapping and Generation
After the cropping and reenactment process, we are ready to fulfil the face swapping process. We first feed the driven pair ($I_{\rm{d}},M_{\rm{d}}$) and the target pair ($I_{\rm{t}},M_{\rm{t}}$) into a mask-guided multi-scale encoder $F_{\phi}$ respectively, obtaining the style codes to represent the texture of each facial region. This step can be described as:
$S_{\rm{t}}=F_{\phi}(I_{\rm{t}},M_{\rm{t}}),\quad S_{\rm{d}}=F_{\phi}(I_{\rm{d}},M_{\rm{d}}),$ (2)
where $S_{\rm{t}}$ and $S_{\rm{d}}$ denote the extracted texture codes of the target and driven face, respectively. The detailed modules of the encoder $F_{\phi}$ are introduced in Sec. III-B. Then, we exchange the texture codes of several facial components of $S_{\rm{t}}$ with those of $S_{\rm{d}}$, resulting in the recomposed texture codes $S_{\rm{swap}}$. The components that must be exchanged from $S_{\rm{d}}$ are: eyebrows, eyes, nose, mouth, lips, face, skin, neck, and ears. Note that due to the entanglement of skin tone and lighting, we split the skin color processing into two steps: (i) transferring the source skin texture including skin tone and lighting from the source to the swapped faces (Sec. III-B); (ii) recoloring the swapped results under the guidance of the target color space (Sec. III-C). The experimental results in Fig. 8 show that our scheme outperforms previous methods, especially on the consistency of source identity.
Next, we feed the recomposed mask $M_{\text{swap}}$ and texture codes ${S_{\text{swap}}}$ into the StyleGAN generator $G_{\theta}$ with a mask-guided style injection module to generate the naive swapped face $\tilde{I}$:
$\tilde{I}=G_{\theta}(M_{\text{swap}},S_{\text{swap}}).$ (3)
The details of the generator $G_{\theta}$ will be introduced in the Sec. III-B. Note that our E4S does not need to train an extra sub-network to deal with the occlusion and can naturally achieve more accurate occlusion recovery than the existing methods FSGAN [13] and FaceShifter [5]. This is because the generator $G_{\theta}$ can fill out the occlusion pixels with the swapped texture features adaptively according to masks.
III-A3 Lighting Transfer and Face Shape Inpainting
Since the disentanglement of texture and shape solely cannot guarantee a lighting-natural swapped result. That is, the lighting of the source contained in the texture information is hard to be disentangled well [52], leading to the lighting leakage to the swapped faces. To tackle this issue and harmonize the lighting results, on the basis of the aforementioned transferring skin texture, we additionally propose a face-recoloring network $B_{\psi}$ to transfer lighting from the target face $I_{\rm{t}}$ to the naive swapped face $\tilde{I}$, which is elaborated in Sec. III-C1. Then, considering that the potential mismatch during the mask recombination process of $M_{\rm{d}}$ and $M_{\rm{t}}$ would bring difficulties for the shape preservation of source face, we propose an inpainting network $P_{\tau}$, which operates on pixel level and maintain the source face shape by the guide of mismatched region masks. we give the detailed description of the inpainting network $P_{\tau}$ in the Sec. III-C2.
III-B Disentangling Shape and Texture
The core of the proposed E4S framework is disentangling the per-region texture and its corresponding shapes. To pursue a better disentanglement of shape and texture as well as high-resolution and high-fidelity generation, we resort to the powerful generative model StyleGAN that can generate images with 1024$\times$1024 resolution naturally. Specifically, we develop a GAN inversion method rather than training StyleGAN from scratch, which also avoids training instability.